OpenStudy (anonymous):

Simplify the following expression x - 4y + 2z - 3x - 3y + 2z = ?

OpenStudy (ageta):

Answer: =−2x−7y+4z

OpenStudy (anonymous):

−2x−7y+4z

OpenStudy (ageta):

answer=-2x-7y+4z

OpenStudy (anonymous):

21

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

:P Sorry

OpenStudy (anonymous):

Let's simplify step-by-step. x−4y+2z−3x−3y+2z =x+−4y+2z+−3x+−3y+2z Combine Like Terms: =x+−4y+2z+−3x+−3y+2z =(x+−3x)+(−4y+−3y)+(2z+2z) =−2x+−7y+4z Answer: =−2x−7y+4z

OpenStudy (ageta):

i am right

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

Do you guys know why its showing all the ? marks?

OpenStudy (anonymous):

aren't you doing that on purpose?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

maybe it shows the marks when you hit the space bar

OpenStudy (anonymous):

probrobly.....well bye

OpenStudy (anonymous):

so its definitely -2x-7y+4z?

OpenStudy (anonymous):

that answer is wrong ageta